Liquidating Partnerships—Capital Deficiency

Nettles, King, and Tanaka are partners sharing income 3:2:1. After the firm's loss from liquidation is distributed, the capital account balances were: Nettles, $24,000 Dr.; King, $87,000 Cr.; and Tanaka, $65,000 Cr.

If Nettles is personally bankrupt and unable to pay any of the $24,000, what will be the amount of cash received by King and Tanaka upon liquidation? If an amount is zero, enter in 0. Use the minus sign to indicate any deficiencies.

Amount of Cash Received
Nettles King Tanaka
Capital balances after realization $ $ $
Distribution of partner deficiency
Capital balances after deficiency distribution $ $ $
Answer #1
Answer

Nettles

King Tanaka
Capital balances after realization -$     24,000 $     87,000 $     65,000
Distribution of partner deficiency $     24,000 -$     16,000 -$       8,000
Capital balances after deficiency distribution $             -   $     71,000 $     57,000
Workings:
Distribution of partner deficiency :
King = 24000*2/3 = $ 16,000
Tanaka = 24000*1/3 = $8,000

Please like